1. What is the basis for the null space of the set "\begin{vmatrix}
 1 &2& 1 \\
 1 &1 & 0\\
-1 & 1 & 0\\
1 & 4 & 1
\end{vmatrix}"
  2. For a given matrix A = "\begin{bmatrix}
 5 & -3 \\
 -6 & 2
\end{bmatrix}" , the matrix P that is diagonalizes A is

(i) P = "\begin{bmatrix}
 1 & 1 \\
 2 & -1
\end{bmatrix}"

(ii) P = "\begin{bmatrix}
 1 & -1 \\
 -2 & 1
\end{bmatrix}"

(iii) P = "\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}"

3.. Suppose T:R2 -> M22 is a linear defined by T(u,v) = "\begin{bmatrix}
 u & v \\
 u & 2u
\end{bmatrix}" then Ker (T) is ...


1
Expert's answer
2021-10-26T17:43:46-0400

1.


"\begin{bmatrix}
 1 & 2 & 1 \\
 1 & 1 & 0 \\
 - 1 & 1 & 0 \\
 1 & 4 & 1 \\
\end{bmatrix}"

Find the reduced row echelon form

"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & 2 & 1 \\
 0 & -1 & -1 \\
 - 1 & 1 & 0 \\
 1 & 4 & 1 \\
\end{bmatrix}"

"R_3=R_3+R_1"


"\begin{bmatrix}
 1 & 2 & 1 \\
 0 & -1 & -1 \\
 0 & 3 & 1 \\
 1 & 4 & 1 \\
\end{bmatrix}"

"R_4=R_4-R_1"


"\begin{bmatrix}
 1 & 2 & 1 \\
 0 & -1 & -1 \\
 0 & 3 & 1 \\
 0 & 2 & 0 \\
\end{bmatrix}"

"R_2=-R_2"


"\begin{bmatrix}
 1 & 2 & 1 \\
 0 & 1 & 1 \\
 0 & 3 & 1 \\
 0 & 2 & 0 \\
\end{bmatrix}"

"R_1=R_1-2R_2"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & 1 \\
 0 & 3 & 1 \\
 0 & 2 & 0 \\
\end{bmatrix}"

"R_3=R_3-3R_2"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & 1 \\
 0 & 0 & -2 \\
 0 & 2 & 0 \\
\end{bmatrix}"

"R_4=R_4-2R_2"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & 1 \\
 0 & 0 & -2 \\
 0 & 0 & -2 \\
\end{bmatrix}"

"R_3=R_3\/(-2)"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & 1 \\
 0 & 0 &1 \\
 0 & 0 & -2 \\
\end{bmatrix}"

"R_1=R_1+R_3"


"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 1 \\
 0 & 0 &1 \\
 0 & 0 & -2 \\
\end{bmatrix}"

"R_2=R_2-R_3"


"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 &1 \\
 0 & 0 & -2 \\
\end{bmatrix}"

"R_4=R_4+2R_3"


"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 &1 \\
 0 & 0 & 0 \\
\end{bmatrix}"

To find the null space, solve the matrix equation


"\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 &1 \\
 0 & 0 & 0 \\
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 \\
 x_3 \\
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
0 \\
0\\
\end{bmatrix}"

Since this system has a unique solution, the null space contains only a zero vector.

The nullity of a matrix is the dimension of the basis for the null space.

Thus, the nullity of the matrix is 0.


The null space is "\begin{bmatrix}
 0 \\
 0 \\
0 \\
\end{bmatrix}," it has no basis.

The nullity of the matrix is "0."


2.


"A=\begin{bmatrix}
 5 &-3 \\
 -6 & 2
\end{bmatrix}"


Find the eigenvalues and eigenvectors


"A-\lambda I=\begin{bmatrix}
 5-\lambda &-3 \\
 -6 & 2-\lambda
\end{bmatrix}"

"\det (A-\lambda I)=\begin{vmatrix}
 5-\lambda &-3 \\
 -6 & 2-\lambda
\end{vmatrix}"

"=(5-\lambda)(2-\lambda)-18=10-7\lambda+\lambda^2-18"

"=\lambda^2-7\lambda-8"

The characteristic equation


"\lambda^2-7\lambda-8=0"

"(\lambda+1)(\lambda-8)=0"

The roots are "\lambda_1=-1, \lambda_2=8."

These are the eigenvalues.

Find the eigenvectors.


"\lambda=-1"

"\begin{bmatrix}
 5-\lambda &-3 \\
 -6 & 2-\lambda
\end{bmatrix}=\begin{bmatrix}
 6 &-3 \\
 -6 & 3
\end{bmatrix}"

"R_2=R_2+R_1"

"\begin{bmatrix}
 6 &-3 \\
 0 & 0
\end{bmatrix}"

"R_1=R_1\/6"

"\begin{bmatrix}
 1 &-1/2 \\
 0 & 0
\end{bmatrix}"

Solve the matrix equation 


"\begin{bmatrix}
 1 &-1/2 \\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0
\end{bmatrix}"

If we take "x_2=t," then "x_1=t\/2."


Thus "\vec x=\begin{bmatrix}
t/2 \\
 t
\end{bmatrix}=\begin{bmatrix}
1 \\
 2
\end{bmatrix}s"

This is the eigenvector.


"\lambda=8"


"\begin{bmatrix}
 5-\lambda &-3 \\
 -6 & 2-\lambda
\end{bmatrix}=\begin{bmatrix}
 -3 &-3 \\
 -6 & -6
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 -3 &-3 \\
 0 & 0
\end{bmatrix}"

"R_1=R_1\/(-3)"


"\begin{bmatrix}
 1 &1 \\
 0 & 0
\end{bmatrix}"

Solve the matrix equation 


"\begin{bmatrix}
 1 &1 \\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0
\end{bmatrix}"

If we take "x_2=t," then "x_1=-t."


Thus "\vec x=\begin{bmatrix}
 -t \\
 t
\end{bmatrix}=\begin{bmatrix}
 1 \\
 -1
\end{bmatrix}s"

This is the eigenvector.


Form the matrix "P," whose column "i" is "i" -th eigenvector


"P=\begin{bmatrix}
 1 & 1 \\
 2 & -1
\end{bmatrix}"

The matrix "P" that is diagonalizes "A" is

(i)


"P=\begin{bmatrix}
 1 & 1 \\
 2 & -1
\end{bmatrix}"

3.


"\begin{matrix}
 1u+0v=0 \\
 0u+1v=0\\
 1u+0v=0 \\
 2u+0v=0\\
\end{matrix}"


"\begin{bmatrix}
 1 & 0 \\
 0 & 1 \\
 1 & 0 \\
 2 & 0 \\
\end{bmatrix}"

Find the reduced row echelon form

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & 0 \\
 0 & 1 \\
 0 & 0 \\
 2 & 0 \\
\end{bmatrix}"

"R_4=R_4-2R_1"


"\begin{bmatrix}
 1 & 0 \\
 0 & 1 \\
 0 & 0 \\
 0 & 0 \\
\end{bmatrix}"

To find the null space, solve the matrix equation


"\begin{bmatrix}
 1 & 0 \\
 0 & 1 \\
 0 & 0 \\
 0 & 0 \\
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 \\

\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
0 \\
0\\
\end{bmatrix}"

Since this system has a unique solution, the null space contains only a zero vector.

The nullity of a matrix is the dimension of the basis for the null space.

Thus, the nullity of the matrix is 0.


The null space is "\begin{bmatrix}
 0 \\
 0 \\

\end{bmatrix}," it has no basis.

The nullity of the matrix is "0." This is the dimension of the kernel of "T."

Ker"(T)" is zero.